QUESTION #1

The method of tree-ring dating gave the following years A.D. for an archaeological excavation site. Assume that the population of x values has an approximately normal distribution.

1,299 1,229 1,215 1,257 1,268 1,316 1,275 1,317 1,275

(a)

Use a calculator with mean and standard deviation keys to find the sample mean year x and sample standard deviation s.

(b)

Find a 90% confidence interval for the mean of all tree-ring dates from this archaeological site.

QUESTION # 2

How much do wild mountain lions weigh? Adult wild mountain lions (18 months or older) captured and released for the first time in the San Andres Mountains gave the following weights (pounds):

68 106 129 125 60 64

Assume that the population of x values has an approximately normal distribution.

(a) Use a calculator with mean and sample standard deviation keys to find the sample mean weight x and sample standard deviation s. (Round your answers to one decimal place.)

x = _____ lb
s = ______ lb


(b) Find a 75% confidence interval for the population average weight μ of all adult mountain lions in the specified region. (Round your answers to one decimal place.)

lower limit     _________ lb
upper limit     _________ lb
Answer #1

Question 1

Values ( X ) Σ ( Xi- X̅ )2
1299 711.1129
1229 1877.7749
1215 3287.1073
1257 235.1101
1268 18.7775
1316 1906.7807
1275 7.1113
1317 1995.1141
1275 7.1113
Total 11451 10046.0001

Part a)

Mean X̅ = Σ Xi / n
X̅ = 11451 / 9 = 1272.3333


Sample Standard deviation SX = √ ( (Xi - X̅ )2 / n - 1 )
SX = √ ( 10046.0001 / 9 -1 ) = 35.4366

Part b)

Confidence Interval
X̅ ± t(α/2, n-1) S/√(n)
t(α/2, n-1) = t(0.1 /2, 9- 1 ) = 1.86
1272.3333 ± t(0.1/2, 9 -1) * 35.4366/√(9)
Lower Limit = 1272.3333 - t(0.1/2, 9 -1) 35.4366/√(9)
Lower Limit = 1250.3626
Upper Limit = 1272.3333 + t(0.1/2, 9 -1) 35.4366/√(9)
Upper Limit = 1294.304
90% Confidence interval is ( 1250.3626 , 1294.304 )

Question 2

Values ( X ) Σ ( Xi- X̅ )2
68 576
106 196
129 1369
125 1089
60 1024
64 784
Total 552 5038

Part a)

Mean X̅ = Σ Xi / n
X̅ = 552 / 6 = 92
Sample Standard deviation SX = √ ( (Xi - X̅ )2 / n - 1 )
SX = √ ( 5038 / 6 -1 ) = 31.7

Part b)

Confidence Interval
X̅ ± t(α/2, n-1) S/√(n)
t(α/2, n-1) = t(0.25 /2, 6- 1 ) = 1.301
92 ± t(0.25/2, 6 -1) * 31.7/√(6)
Lower Limit = 92 - t(0.25/2, 6 -1) 31.7/√(6)
Lower Limit = 75.2
Upper Limit = 92 + t(0.25/2, 6 -1) 31.7/√(6)
Upper Limit = 108.8
75% Confidence interval is ( 75.2 , 108.8 )